Some of the information found on a detail inventory card for Novak Inc. for the first month of operations is as follows. If the perpetual inventory record is kept in dollars, and costs are computed at the time of each withdrawal, would the amounts shown as ending inventory in (1)FIFO, (2)LIFO, and (3)Average Cost above be the same? What amount would be shown as ending inventory? (Round average cost per unit to 4 decimal places, e.g. 2.7621 and final answers to 0 decimal places, e.g. 6,548.)

Received

Date

No. of Units

Unit Cost

Issued,
No. of Units

Balance,
No. of Units

January     2 1,700 $5.19 1,700
7 1,200 500
10 1,100 5.54 1,600
13 1,000 600
18 1,500 5.71 800 1,300
20 1,100 200
23 1,800 5.88 2,000
26 1,300 700
28 2,100 6.06 2,800
31 1,800

1,000
Answer #1

LIFO METHOD

Date Recived Issued Balance inventory
January 2 1,700×5.19=8,823 1,700×5.19=8,823
January 7 1,200×5.19=6,228 500×5.19=2,595
January 10 1,100×5.54=6,094

500×5.19=2,595

1,100×5.54=6,094

January 13 1,000×5.54=5,540

500×5.19=2,595

100×5.54=554

January 18 1,500×5.71=8,565 800×5.71=4,568

500×5.19=2,595

100×5.54=554

700×5.71=3,997

January 20

700×5.54=3,997

100×5.54=554

300×5.19=1,557

200×5.19=1,038
January 23 1,800×5.88=10,584 200×5.19=1,038
January 26 1,300×5.88=7,644

200×5.19=1,038

500×5.88=2,940

January 28 2,100×6.06=12,726

200×5.19=1,039

500×5.88=2,940

2,100×6.06=12,726

January 31 1,800×6.06=10,908

200×5.19=1,038

500×5.88=2,940

300×6.06=1,818

Total Ending inventory = 1,038+2,940+1,818= $5,796

FIFO METHOD

Date Received Issued Balance inventory
January 2 1,700×5.19=8,823 1,700×5.19=8,823
7 1,200×5.19=6,228 500×5.18=2,595
10 1,100×5.54=6,094

500×5.19=2,595

1,100×5.54=6,094

13

500×5.19=2,595

500×5.54=2,770

600×5.54=3,324
18 1,500×5.71=8,565

600×5.54=3,324

200×5.71=1,142

1,300×5.71=7,423
20 1,100×5.71=6,281 200×5.71=1,142
23 1,800×5.58=10,584

200×5.71=1,142

1,800×5.88=10,584

26

200×5.71=1,142

1,100×5.88=6,468

700×5.88=4,116
28 2,100×6.06=12,726

700×5.88=4,116

2,100×6.06=12,726

31

700×5.88=4,116

1,100×6.06=6,666

1000×6.06=6,060

Total cost of ending inventory= $6,060.

Weighted average cost method

Date Received Issued Balance inventory
January 2 1,700×5.19=8,823 1,700×5.19=8,823
7 1,200×5.19=6,228 500×5.19=2,595
10 1,100×5.54=6,094 1,600×5.43=8,689
13 1000×5.43=5,430 600×5.43=3,259
18 1,500×5.71=8,565 800×5.63=4,505 1,300×5.63=7,319
20 1,100×5.63=6,193 200×5.63=1,126
23 1,800×5.85=10,584 2,000×5.85=11,710
26 1,300×5.85=7,611 700×5.85=4,098
28 2,100×6.06=12,726 2,800×6.008=16,824
31 1,800×6.008=10,816 1000×6.008=6,008

Total cost of ending inventory= $6,008.

____×____

Comment if you have any doubt

Keep supporting.